OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates. (-7, 2pi/3)? I think its (7/2, -7sqrt3/2) ??

OpenStudy (amistre64):

given a polar (r,a) the cartesian is: (r cosa , r sina)

OpenStudy (amistre64):

-7 is an odd construct tho ... valid; but odd turning 2pi/3 and backing up by 7 is the same: (2-3)pi/3 = -pi/3 60 under x

OpenStudy (amistre64):

pfft, pictures off

OpenStudy (amistre64):

\[\frac{2\cancel{\pi}~~180}{3~~~~\cancel{\pi}}\] \[2(60)=120\] \[120-180=-60^o\]

OpenStudy (anonymous):

..... i have no clue what you just did :(

OpenStudy (dumbcow):

you are correct

OpenStudy (amistre64):

i was trying to find the appropriate angle is all.... (-7,rad) is an odd construct; you have to move backwards to get to the terminal point is all

OpenStudy (dumbcow):

yeah normally r is positive...a neg "r" adds pi to the angle

OpenStudy (amistre64):

its best to use (7,a-pi)

OpenStudy (amistre64):

add or subtract:)

OpenStudy (dumbcow):

right either 1 works

OpenStudy (amistre64):

you end up in the 4th quad as a result which means your x and y are going to have negative values ... so unless theres a typo your results were in error

OpenStudy (amistre64):

opps, im in error now

OpenStudy (amistre64):

+x,-y in Q4

OpenStudy (anonymous):

im so confused guys lol..

OpenStudy (amistre64):

the outcome is still valid for: x = r cosa, y=r sina

OpenStudy (amistre64):

-7 cos(2pi/3) = 7/2 -7 sin(2pi/3) = -7sqrt(3)/2

OpenStudy (anonymous):

okay thankyou lol(:

OpenStudy (amistre64):

:) good luck